Can someone help me solve this problem?
At 1:00 p.m., a car leaves a city and travels north at a rate of 55 mi/h. An hour later, a second car leaves the city and travels south at a rate of 60 mi/hr. At what time will the two cars be 285 miles apart?
Thanks, Lyn

Answer by stanbon(48544)

At 1:00 p.m., a car leaves a city and travels north at a rate of 55 mi/h. An hour later, a second car leaves the city and travels south at a rate of 60 mi/hr. At what time will the two cars be 285 miles apart?
---------------
North car DATA:
rate=55 mph ; time= x hrs. ; distance= 55x mi.
---------------
South car DATA:
rate= 60 mph ; time= x-1 hrs ; distance = 60(x-1)=60x-60 miles
------------------
EQUATION:
south distance+ north distace = 285 miles
60x-60+55x=285
115x=345
x=3 hrs
1:00p.m.+3 hrs. = 4:00p.m.
Cheers,
Stan H.